A fall in price of from 1 shs 0.50 per bushel to 9.20 per bushel, the quantity of the demand from 18800 to 21200 per bushells. What is the price elasticity of demand?




1
Expert's answer
2022-01-21T09:34:04-0500

Solution:

Price elasticity of demand (PEd) = "=\frac{\%\;change\; in\; quantity\; demanded}{\%\; change\; in\; price}"

Q1 = 18800        Q2 = 21200

P1 = 1                  P2 = 0.50

 

PED = "=\frac{Q_{2} -Q_{1}}{(Q_{2}+Q_{1})/2 } \div \frac{P_{2} -P_{1}}{(P_{2}+P_{1})/2 }"


PED = "\frac{21200 -18800}{(21200+18800)/2 } \div \frac{21200 -18800}{(21200+18800)/2 }"


PED ="\frac{2400}{20000} \div\frac{-0.5}{0.75} = \frac{0.12}{-0.67} = -0.18"


The price elasticity of demand = -0.18